# Multi-Scale Grouped Prototypes for Interpretable Semantic Segmentation

### v1

@@ 1 Introduction
In the last few years, deep learning-based semantic segmentation has seen rapid adoption in numerous fields, from industrial use cases such as autonomous driving [22, 74, 81] to the environmental sciences [65, 5, 48]. This expansion was driven by its increasing performance on a multitude of tasks and benchmarks, often coinciding with an increase in model complexity [89, 64, 11, 86]. This favors un-interpretable black-box models, to the detriment of their explainability.
@@ 2 Related works
Among the few existing methods tackling this problem, post-hoc approaches extending Grad-CAM to semantic segmentation have been explored [83, 35]. By-design approaches can provide explainable results by leveraging symbolic language [71], through the use of a semantic bottleneck [56] or by exploiting the attention mechanism [28]. Prototypical parts learning was proposed recently in ProtoSeg [70] for interpretable semantic segmentation, extending the classification-based method proposed in [8]. In this work, we aim to investigate multi-scale prototype learning to break the performance/interpretability trade-off present in ProtoSeg. This is achieved by explicitly leveraging the multi-scale nature of semantic segmentation representations (via scale-specific prototypes) and by encouraging sparsity (via sparse grouping across prototypes and scales).
@@ 3.1 Multi-scale prototype learning
$w_{h_{\text{proto}}}^{(c,\operatorname*{m})}=\left\{\begin{aligned} &1&&\text{if}\ \mathbf{p}_{s,\operatorname*{m}}\in P_{c}\\ &-0.5&&\text{otherwise}\end{aligned}\right.$
@@ 3.3 Multi-stage training procedure
$L_{\text{ent}}=\frac{1}{C\cdot N}\sum_{c\in\mathcal{C}}\sum_{\operatorname*{n}=1}^{N_{c}}\sum_{\operatorname*{m}=1}^{M}-w_{g,c}^{(\operatorname*{n},\operatorname*{m})}\log(w_{g,c}^{(\operatorname*{n},\operatorname*{m})})$ (8)
Then, in the second step, we fine-tune also the linear layer $h_{\text{group}}$. To enforce more sparsity, following the training protocol from [8], we apply an L1-norm loss term on $h_{\text{group}}$ solely on the weights $w_{h_{\text{group}}}^{c,\operatorname*{n}}$ where the group $\mathbf{g}_{\operatorname*{n}}$ is not in the set of groups $G_{c}$ assigned to class $c$. The total loss term for the grouping mechanism (Figure 2: Stage 2) becomes:
$L_{\text{proto}}=L_{\text{CE}}+\lambda_{\text{ent}}\cdot L_{\text{ent}}+\lambda_{\text{L1}}\cdot\sum_{c\in\mathcal{C}}\sum_{\operatorname*{n}:\mathbf{g}_{\operatorname*{n}}\notin G_{c}}|w_{h_{\text{group}}}^{(c,\operatorname*{n})}|$ (9)
@@ 4.1 Experimental setup
In all the experiments presented in the section 4.2, we use DeepLabv2 [9] with ResNet-101 [33] pre-trained on ImageNet as the backbone. For the multi-scale prototype learning, we assign $M=3$ scale-dependent prototypes per scale to each class and $S=4$ scales, so $12$ prototypes per class in total. Moreover, we set the number of groups per class to $N_{c}=3$ for the grouping mechanism. We evaluate ScaleProtoSeg on (i) Pascal VOC 2012 [21] (made of $1464$ train, $1449$ validation, and $1446$ test images with $21$ classes; the Pascal VOC training set is extended to $10582$ images following [29]), (ii) Cityscapes [14] (composed of $2975$ train, $500$ validation, and $1525$ test images of street scenes, with $19$ classes) and (iii) ADE20K (a scene-parsing dataset with 150 fine-grained semantic classes split in $\sim 20000$ training and $2000$ validation images; for the training of ProtoSeg [70] on ADE20K we extend the number of prototypes to $12$ for direct comparison with ScaleProtoSeg). A detailed description of the experimental setup is available in the supplementary materials in Section 6.
@@ 4.2 Results and discussion
Table 2: Methods performance at different training steps based on the results from [70] and our best ScaleProtoSeg run. The number of prototypes indicated in the project step (during Stage 1) is after the duplicate removal. The number of prototypes in the grouped step (during Stage 2) corresponds to a threshold of $0.05$ on the group matrices.
@@
In terms of interpretability, the groups can be first represented through specific visualization for each sample, as shown in Figure 4, where the images on the graph display the prototype activations for an input sample. Those activations are grouped per scale and provide local explanations for the model. Moreover, the edges in those visualizations are static for all samples assigned to the class car in Cityscapes, and so provide global explanations on the model behavior. In Figure 4, we can easily observe, despite the large number of prototypes compared to the model mean of $3.39$, that different patterns emerge for each group. The first group focuses mainly on a prototype in Scale 4 activated on the bottom part of the cars and another similar at Scale 2. The second group focuses on a prototype at Scale 1, which activates the top of the cars with another one similar at Scale 3. Lastly, the third group focuses on multiple parts of the cars and, will activate on the main body of them. Groups $1$ and $2$, by combining similar prototypes across two scales, show transparently the use of multi-scale information in the model. Figure 4 demonstrates that the grouping mechanism in our method can be represented for both local and global interpretability. The example of group assignments for the class car illustrates the identified parts in the group visualization. Interestingly, the second group seems more activated on cars further in the depth field, as its main activated prototype is from Scale 1 with a smaller field of view. This clearly illustrates the role of scale in the identification of prototypical parts.
@@ 6 Details on experimental set-up
Firstly, across all training stages, the prototypes are the same size as the output feature maps of the non-concatenated ASPP in DeepLabv2 [9]: $D=64$. For both multi-scale prototype training and the grouping mechanism we leverage augmentation techniques, such as random horizontal flipping, cropping, and scaling images by a factor between $[0,1.5]$ for Cityscapes [14] and Pascal VOC [21], and $[0,2]$ for ADE20K [100]. The batch size used in all the training stages is $10$ and we select the Adam optimizer with a weight decay of $5e^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$ for our experiments. Moreover, as the batch size in our experiments is limited, we freeze the batch normalization parameters of the ResNet-101 backbone due to its impact on performance.
@@ 8 Extension to large dataset
In order to test our method in a use case close to real-world applications we extend our evaluation to a large benchmark: COCO-Stuff [7], containing 182 classes with 118k training and 5k validation samples. We leverage a similar setup to ADE20K for the model hyperparameters, except that in the multi-scale prototype training stage, we iterate for a total of 110k iterations: 6k warm-up steps, 100k joint steps, and 4k finetuning steps. Moreover, the iterations are done directly on a batch of size 10. We also use multi-scale inputs with scales $\{0.5,0.75,1\}$ and a training resolution of $321\times 321$, while testing at the native resolution. The objective is to align to the set-up for the DeepLabv2 baseline provided at this repository: github.com/kazuto1011/deeplab-pytorch. The results for our method: ScaleProtoSeg are presented in Table 6, and showcase that our method despite lower performance stays competitive with DeepLabv2 on a larger dataset while providing interpretability. Moreover, ProtoSeg trained in a similar set-up has a performance of 32.97 mIoU without pruning for one run, and so ScaleProtoSeg slightly outperforms it.
@@ 13 Group Representations and Assignments
Figure 11: ScaleProtoSeg provides an interpretation of the resulted segmentation through groups of prototypes. For an example of the class person on Cityscapes, $2$ prototypes per scale (whose activations are displayed at the top of the figure) are used by the model across the $3$ learned groups shown at the bottom right. For this class, groups can correspond to the feet, the main body or the head of the person.

### v2

@@ 1 Introduction
In the last few years, deep learning-based semantic segmentation has seen rapid adoption in numerous fields, from industrial use cases such as autonomous driving [22, 74, 81] to the environmental sciences [65, 5, 48]. This expansion was driven by its increasing performance on a multitude of tasks and benchmarks, often coinciding with an increase in model complexity [89, 64, 11, 86], which favors un-interpretable black-box models, to the detriment of their explainability.
@@ 2 Related works
Among the few existing methods tackling this problem, extensions of Grad-CAM to semantic segmentation have been explored [83, 35] for post-hoc methods. By-design approaches can provide explainable results by leveraging symbolic language [71], through the use of a semantic bottleneck [56] or by exploiting the attention mechanism [28]. Prototypical parts learning was proposed recently in ProtoSeg [70] for interpretable semantic segmentation, extending the classification-based method proposed in [8]. In this work, we aim to investigate multi-scale prototype learning to break the performance/interpretability trade-off present in ProtoSeg. This is achieved by explicitly leveraging the multi-scale nature of semantic segmentation representations (via scale-specific prototypes) and by encouraging sparsity (via sparse grouping across prototypes and scales).
@@ 3.1 Multi-scale prototype learning
$w_{h_{\text{proto}}}^{(c,\operatorname*{m})}=\left\{\begin{aligned} &1&&\text{if}\ \mathbf{p}_{\operatorname*{m}}\in P_{c}\\ &-0.5&&\text{otherwise}\end{aligned}\right.$
@@ 3.3 Multi-stage training procedure
$L_{\text{ent}}=\frac{1}{C\cdot N_{c}}\sum_{c\in\mathcal{C}}\sum_{\operatorname*{n}=1}^{N_{c}}\sum_{\operatorname*{m}=1}^{M}-w_{g,c}^{(\operatorname*{n},\operatorname*{m})}\log(w_{g,c}^{(\operatorname*{n},\operatorname*{m})})$ (8)
Then, in the second step, we fine-tune also the linear layer $h_{\text{group}}$. To enforce more sparsity, following the training protocol from [8], we apply an L1-norm loss term on $h_{\text{group}}$ solely on the weights $w_{h_{\text{group}}}^{c,\operatorname*{n}}$ where the group $g_{\operatorname*{n}}$ is not in the set of groups $G_{c}$ assigned to class $c$. The total loss term for the grouping mechanism (Figure 2: Stage 2) becomes:
$L_{\text{proto}}=L_{\text{CE}}+\lambda_{\text{ent}}\cdot L_{\text{ent}}+\lambda_{\text{L1}}\cdot\sum_{c\in\mathcal{C}}\sum_{\operatorname*{n}:g_{\operatorname*{n}}\notin G_{c}}|w_{h_{\text{group}}}^{(c,\operatorname*{n})}|$ (9)
@@ 4.1 Experimental setup
In all the experiments presented in the section 4.2, we use DeepLabv2 [9] with ResNet-101 [33] pre-trained on ImageNet as the backbone. For the multi-scale prototype learning, we assign $M=3$ scale-dependent prototypes per scale to each class and $S=4$ scales, so $12$ prototypes per class in total. Moreover, we set the number of groups per class to $N_{c}=3$ for the grouping mechanism. We evaluate ScaleProtoSeg on (i) Pascal VOC 2012 [21] (made of $1464$ train, $1449$ validation, and $1446$ test images with $21$ classes; the Pascal VOC training set is extended to $10582$ images following [29]), (ii) Cityscapes [14] (composed of $2975$ train, $500$ validation, and $1525$ test images of street scenes, with $19$ classes) and (iii) ADE20K [100] (a scene-parsing dataset with 150 fine-grained semantic classes split in $\sim 20000$ training and $2000$ validation images; for the training of ProtoSeg [70] on ADE20K we extend the number of prototypes to $12$ for direct comparison with ScaleProtoSeg). A detailed description of the experimental setup is available in the supplementary materials in Section 6.
@@ 4.2 Results and discussion
Table 2: Methods performance at different training steps based on the results from [70] and our best ScaleProtoSeg run. The number of prototypes indicated in the project step (during Stage 1) is after the removal of the duplicates. The number of prototypes in the grouped step (during Stage 2) corresponds to a threshold of $0.05$ on the group matrices.
@@
In terms of interpretability, the groups can be first represented through specific visualization for each sample, as shown in Figure 4, where the images on the graph display the prototype activations for an input sample. Those activations are grouped per scale and provide local explanations for the model. Moreover, the edges in those visualizations are static for all samples assigned to the class car in Cityscapes, and so provide global explanations on the model behavior. In Figure 4, we can easily observe, despite the large number of prototypes compared to the model mean of $3.39$, that different patterns emerge for each group. The first group focuses mainly on a prototype in Scale 4 activated on the bottom part of the cars and another similar at Scale 2. The second group focuses on a prototype at Scale 1, which activates the top of the cars with another one similar at Scale 3. Lastly, the third group focuses on multiple parts of the cars and, will activate on the main body of them. Groups $1$ and $2$, by combining similar prototypes across two scales, show transparently the use of multi-scale information in the model. Figure 4 demonstrates that the grouping mechanism in our method can be representative of both local and global interpretability. The example of group assignments for the class car illustrates the identified parts in the group visualization. Interestingly, the second group seems more activated on cars further in the depth field, as its main activated prototype is from Scale 1 with a smaller field of view. This clearly illustrates the role of scale in the identification of prototypical parts.
@@ 6 Details on experimental set-up
Firstly, across all training stages, the prototypes are the same size as the output feature maps of the non-concatenated ASPP in DeepLabv2 [9]: $D=64$. For both multi-scale prototype training and the grouping mechanism we leverage augmentation techniques, such as random horizontal flipping, cropping, and scaling images by a factor between $[0.5,1.5]$ for Cityscapes [14] and Pascal VOC [21], and $[0.5,2]$ for ADE20K [100]. The batch size used in all the training stages is $10$ and we select the Adam optimizer with a weight decay of $5e^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$ for our experiments. Moreover, as the batch size in our experiments is limited, we freeze the batch normalization parameters of the ResNet-101 backbone due to its impact on performance.
@@ 8 Extension to large dataset
In order to test our method in a use case close to real-world applications we extend our evaluation to a large benchmark: COCO-Stuff [7], containing 182 classes among 11 are not present in the dataset with 118k training and 5k validation samples. We leverage a similar setup to ADE20K for the model hyperparameters, except that in the multi-scale prototype training stage, we iterate for a total of 110k iterations: 6000 warm-up steps, 100k joint steps, and 4000 finetuning steps. Moreover, the iterations are done directly on a batch of size 10. We also use multi-scale inputs with scales $\{0.5,0.75,1\}$ and a training resolution of $321\times 321$, while testing at the native resolution. The objective is to align to the set-up for the DeepLabv2 baseline provided at this repository: github.com/kazuto1011/deeplab-pytorch. The results for our method: ScaleProtoSeg are presented in Table 6, and showcase that our method despite lower performance stays competitive with DeepLabv2 on a larger dataset while providing interpretability. Moreover, ProtoSeg trained in a similar setup has a performance of 32.97 mIoU without pruning for one run, so ScaleProtoSeg slightly outperforms it.
@@ 13 Group Representations and Assignments
Figure 11: ScaleProtoSeg provides an interpretation of the resulting segmentation through groups of prototypes. For an example of the class person on Cityscapes, $2$ prototypes per scale (whose activations are displayed at the top of the figure) are used by the model across the $3$ learned groups shown at the bottom right. For this class, groups can correspond to the feet, the main body or the head of the person.